78% of the area under the standard normal curve is between -z and z.

how this can be answered?
Answer #1

p(z<z) -p (Z<-2)=0.78 p(z<z)-p (Z>Z)=0.78 2*p(2<z)=1.78 p(2<z)=0.89 20.89=1.2265 and -z=-1.2265